OpenStudy (anonymous):

log2(x+5)=4+log2(x-1) Solve for x.

OpenStudy (anonymous):

log2(x - 4) + log2(x) = 5 : x > 4 log2[x(x-4)] = 5 x(x-4) = 2^5 = 32 x² - 4x - 32 = 0 (x+4)(x-8) = 0 x = -4, 8 : but since x > 4, discard the extraneous x = -4 result x = 8

OpenStudy (mertsj):

Subtract log2(x-1) from both sides.

OpenStudy (anonymous):

that is not an option the options are a. {-2/5} b. { 2/5} c. {7/5} d. {-7/5}

OpenStudy (anonymous):

@Mertsj I got log2(x+5)-log2(x-1)=4

OpenStudy (mertsj):

Now use the law that says that \[\log_{b}m-\log_{b} n=\log_{b}\frac{m}{n} \]

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

@Mertsj I dont know anything about a law? But what about the 4?

OpenStudy (mertsj):

\[\log_{2} \frac{x+5}{x-1}=4\]

OpenStudy (mertsj):

Now put it into exponential form. What is the base?

OpenStudy (anonymous):

you lost me with expo form but the base is 2 right?

OpenStudy (mertsj):

And the log is the exponent so what is the exponent?

OpenStudy (mertsj):

So you don't know the laws of logs and you don't know the exponential form of a log equation so why are you doing this problem? What are you supposed to be studying?

OpenStudy (anonymous):

I know the exponential form of a log but I dont know how to do it for this equation. So do I move the 4 over or not?

OpenStudy (mertsj):

The log is the exponent. So what is the exponent?

OpenStudy (anonymous):

I don't know is it 1?

OpenStudy (mertsj):